OpenStudy (anonymous):

4. How many ways can 11 Redcoat basketball players and Coach Woodburn (12 people) be arranged along the sidelines: a) with no restrictions. b) if Coach Woodburn must be at one end of the bench? c) if Travis and Pat must stand together? d) if Travis, Pat, and Woody must stand together? e) if Travis and Pat must be kept separated? Still checking answers: as follows

OpenStudy (anonymous):

a) 12i = 479 001 600 b) 399 16800 c) 79833600 d) " " e)399 16 8000

OpenStudy (anonymous):

my problem is whether i should have used nCr methods for these..

OpenStudy (anonymous):

i assume the first one you mean \(12!\) right?

OpenStudy (anonymous):

@satellite73 then again..you know that's the problem huh. hahha. yes!

OpenStudy (anonymous):

oops. yeah i didn't mean imaginary numbers lol.

OpenStudy (anonymous):

ok now let us dispense with formulas and think only using the counting principle notice you used no formula for the first one, 12 choices, then 11, then 10, ...

OpenStudy (anonymous):

okay... yeah i just did 12 x 11 x ....etc

OpenStudy (anonymous):

in the second one the coach is at either and of the bench how many way if he is in the first seat?

OpenStudy (anonymous):

i said two ways...front or back so i just calculated 11! then multiplied that by 2

OpenStudy (anonymous):

that is what i would do and notice we appealed to no formula

OpenStudy (anonymous):

true....so none of these use a formula at all?? just the counting/multiplication?

OpenStudy (anonymous):

unfortunately i do not get the same number you wrote

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2*11!

OpenStudy (anonymous):

i think maybe you got b and c backwards

OpenStudy (anonymous):

oops that is what i got i typed it in differently than i had on my sheet

OpenStudy (anonymous):

but i got the same for c

OpenStudy (anonymous):

so you were thinking clearly for C as well. treat travis and pat as one person, then multiply by 2

OpenStudy (anonymous):

yep... and for D 3 as 1 person ... 6*10!

OpenStudy (anonymous):

so answer for b and c are both \(2\times 11!\)

OpenStudy (anonymous):

right and why? because \(3\times 2=6\) possible arrangements for the three of them and then \(10!\) ways to put them in the slots so once again it is counting principle all the way

OpenStudy (anonymous):

Yes that is what I got..cept I did an amateur tree diagram :P

OpenStudy (anonymous):

w/e it worked

OpenStudy (anonymous):

basically what i DONT get is why this question all uses counting and not nCr formulas.....uuurggh this is frustrating

OpenStudy (anonymous):

not sure what that is, but ok now for the last one i do not know a snap way to do it

OpenStudy (anonymous):

what i am trying to suggest is that it will be easier for you (and you seem to have a decent handle on this) if you stop thinking along the lines of "what formula should i use" and keep thinking "what is going on and how to i count this?"

OpenStudy (anonymous):

okay I did 11! leaving one or the other out...than assumably theres 12 places to put the other person minus 2 because there are two places they can't be... so I multiplied 11! by 10

OpenStudy (anonymous):

you use the formula you need when you need them, but don't get married to any formula. each one is based on the counting principle

OpenStudy (anonymous):

right..makes sense

OpenStudy (anonymous):

really it is true and in fact one should never even compute the formula as written suppose i want \(\dbinom{10}{7}\) what do i do? first i think " choosing 7 from 10 is identical to choosing 3 from 10" then i think 10 choices for first slot, 9 for second, 8 for third, all divided by number of ways to arrange these three then i write \[\dbinom{10}{7}=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

thats true. but all too often teachers want you to use formulas :P

OpenStudy (anonymous):

math teachers suck in general, but in an on line class who cares in an on line class i would compute 10 choose 3 like this http://www.wolframalpha.com/input/?i=10+choose+3

OpenStudy (anonymous):

i think we need to be careful with this one e) if Travis and Pat must be kept separated?

OpenStudy (anonymous):

yeah exactly..well i've had to resort to places like this (not in a bad way) because the teacher doesn't answer questions...i understand its summer but i'm a paranoid nerd who needs constant help lol

OpenStudy (anonymous):

okay..assumably my answer is wrong

OpenStudy (anonymous):

oh i don't know, i didn't do it

OpenStudy (anonymous):

i just did 10x11!

OpenStudy (anonymous):

but i think the number of possibilities is different depending on whether travis is in the first bench or in the middle

OpenStudy (anonymous):

..urggggg

OpenStudy (anonymous):

if travis in in the first bench , then in the second bench there are 10 choices and then \(10!\) from there on in

OpenStudy (anonymous):

likewise if travis is in the last bench

OpenStudy (anonymous):

right, makes sense..

OpenStudy (anonymous):

but it travis is in the second bench, then there are 10 choices in the first bench, 9 choices for the third, and then 9! from there on in so i think you have some work to do

OpenStudy (anonymous):

it would appear so.....

OpenStudy (anonymous):

but it is not too bad

OpenStudy (anonymous):

okay so wait

OpenStudy (anonymous):

okay no I don't get it.. never mind -.-

OpenStudy (anonymous):

if we label the benches 1 through 12 and put travis in a specific bench, we can compute easily enough bench 1 or 12 \(10\times 10!\)

OpenStudy (anonymous):

10 x 10! makes more sense... i've resorted to drawing out benches which is relatively useless

OpenStudy (anonymous):

then in the other boxes i think it is a different calculation

OpenStudy (anonymous):

i am sorry i am so confused on this one

OpenStudy (anonymous):

box 2 would be 10 choices for bench 1 and then 9 choices for bench 3 and then 9! choices for the remaining 9 benches does that look rigth?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lets put travis in bench 2 _ T _ _ _ _ _ _ _ _ _ _

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

how many choice for bench one?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

that is what i get how many for bench 3?

OpenStudy (anonymous):

also 10

OpenStudy (anonymous):

..i think

OpenStudy (anonymous):

i think maybe 9

OpenStudy (anonymous):

we have travis seated we have one other person seated and we cannot seat Pat

OpenStudy (anonymous):

couldn't the options for 1 and 3 be interchangeable

OpenStudy (anonymous):

yes, but i am just counting as usual ten choices for first seat, anyone but travis or pat one choice for second seat, travis 9 choices for third seat not the guy in the first seat, not travis, not pat

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so thats where you got 10 x 10!

OpenStudy (anonymous):

that is where i got \(10\times 10!\) from in the first case , if travis is in seat one

OpenStudy (anonymous):

wait but this will be different T _ _ _ _ _ _ _ _ _ _ _

OpenStudy (anonymous):

yes, that will be the \(10\times 10!\) i think ten choices for bench 2 (everyone but travis or pat) then the remaining ten seats with no restrictions so \(10\times 10!\)

OpenStudy (anonymous):

so they both are the same..?

OpenStudy (anonymous):

for this one _ T _ _ _ _ _ _ _ _ _ _ i get \[10\times 9\times 9!=9\times 10!\]

OpenStudy (anonymous):

okay so...... 9 x 10 ! + 10 x 10!

OpenStudy (anonymous):

2 min..grilled cheese is calling me

OpenStudy (anonymous):

here is my reasoning bench one not travis, not pat, ten choices bench two travis one choice bench 3 not travis, not pat, not guy if first bench, 9 choices other 9 benches, no restrictions, 9! choices

OpenStudy (anonymous):

mmm lunch

OpenStudy (anonymous):

oh hold one what a much easier way to do this!! now i feel like a moron take your answer to the first one "a) with no restrictions." and subtract your answer to the third one "c) if Travis and Pat must stand together?" sorry i did not see this right away now you are done

OpenStudy (anonymous):

seriously :o

OpenStudy (anonymous):

i got.... 39916800. ohdear. that was my original answer :P

OpenStudy (anonymous):

thank you sooo much tho!

OpenStudy (anonymous):

well i do feel silly for not seeing this right away

OpenStudy (anonymous):

i'm going to give up on this assignment now and pray the rest of the questions are right. getting quite sick of it. thank you so much for all your help!

OpenStudy (anonymous):

you are quite welcome